calculate the number of moles of sodium oxide that should be produced when 26.50 grams of sodium metal reacts with oxygen gas


1
Expert's answer
2021-04-14T06:10:52-0400

4Na+O2=2Na2O

n(Na)=m(Na)/M(Na)= 26.50g / 23g/mol = 1.152 mol

n(Na2O)= n(Na)/2 = 1.152mol/2= 0.576 mol


Answer: 0.576 mol Na2O